OpenStudy (anonymous):

Solve x^2 + 6x + 7 = 0.

myininaya (myininaya):

\[x^2+6x=-7 \] \[x^2+6x+9=-7+9\] \[(x+3)^2=2\] \[x+3=\pm \sqrt{2}\] \[x=-3 \pm \sqrt{2}\]